# Learning to Evolve: Bayesian-Guided Continual Knowledge Graph Embedding

### v1

@@
Abstract
Since knowledge graphs (KG) will continue to evolve in real scenarios, traditional KGE models are only suitable for static knowledge graphs. Therefore, continual knowledge graph embedding (CKGE) has attracted the attention of researchers. Currently, a key challenge facing CKGE is that the model is prone to ”catastrophic forgetting”, resulting in the loss of previously learned knowledge. In order to effectively alleviate this problem, we propose a new CKGE model BAKE. First, we note that the Bayesian posterior update principle provides a natural continual learning strategy that is insensitive to data order and can theoretically effectively resist the forgetting of previous knowledge during data evolution. Different from the existing CKGE method, BAKE regards each batch of new data as a Bayesian update of the model prior. Under this framework, as long as the posterior distribution of the model is maintained, the model can better preserve the knowledge of early snapshots even after evolving through multiple time snapshots. Secondly, we propose a continual clustering method for CKGE, which further directly combats knowledge forgetting by constraining the evolution difference (or change amplitude) between new and old knowledge between different snapshots. We conduct extensive experiments on BAKE on multiple datasets, and the results show that BAKE significantly outperforms existing baseline models.
Introduction
@@
Knowledge graphs (KGs)(Liang et al. 2024b)(Ji et al. 2021) have become indispensable tools for representing structured information, supporting a wide range of knowledge‑driven applications such as LLM based question answering(Pan et al. 2024)(Xu et al. 2025) and recommender systems(Zhao et al. 2024)(Yuan et al. 2025b). Knowledge graph embedding (KGE) aims to encode entities and relations in a KG as continual vectors, and the quality of these embeddings directly affects downstream performance. Classic KGE models (Bordes et al. 2013)(Li et al. 2023)(Sun et al. 2019)(Xiao, He, and Cao 2023)(Long et al. 2024)(Li et al. 2025c) perform excellently on static KGs, yet real‑world graphs usually evolve continuously. For example, the YAGO knowledge base evolved from YAGO3(Mahdisoltani, Biega, and Suchanek 2013) to YAGO4.5(Suchanek et al. 2024), with the number of entities soaring from about 4.6 million to 49 million and facts (triples) from 40 million to 132 million. For such continually evolving knowledge graphs (CKGs), existing KGE models must retrain the entire graph after each update—a computationally prohibitive and impractical strategy for large‑scale graphs like YAGO. Hence, this work focuses on CKGE models capable of learning CKGs efficiently.
Figure 1: TKG typically employs triplets combined with timestamps to explicitly annotate temporal information on all relational edges, exhibiting a strong temporal dimension. In contrast, CKG decomposes real KGs into a series of time unmarked evolutionary snapshots, focusing on continuously updating and transferring learned knowledge as snapshots evolve.
Recent CKGE studies optimize along three axes: training structure, parameter efficiency, and regularization and masking mechanisms. Structure: IncDE(Liu et al. 2024a) employs a hierarchical strategy with incremental distillation to exploit graph structure. Parameter efficiency: FastKGE(Liu et al. 2024b) reduces trainable parameters via incremental low‑rank adapters, whereas ETT‑CKGE(Zhu et al. 2025) leverages task‑driven tokens for efficient knowledge transfer. Regularization and masking model: existing approaches include lifelong learning based on masked auto‑encoders(Cui et al. 2023), energy‑based alignment of old and new knowledge(Cao 2024), a bio‑inspired dual‑mask mechanism(Song et al. 2024), and flexible regularization grounded on the Fisher information matrix(Zhu et al. 2024). However, these methods largely rely on heuristic regularizers or masks to passively relieve forgetting and have not yet actively learned from the perspective of continuous data evolution by letting the prior guide the posterior.
@@
As illustrated in Figure.1, CKGs differ from temporal KGs(Zhang et al. 2025)(Wang et al. 2023), whose time‑stamped relations explicitly model temporal dependencies. CKGs instead focus on structural increments, such as newly emerging entities or relations(Cui et al. 2023). We observe that, because a CKG lacks explicit temporal semantics, its evolution naturally aligns with the sequential update mechanism of the Bayesian framework. According to Bayes’ rule(Lee et al. 2024)(Bonnet et al. 2025), we treat the posterior distribution of snapshot $t$ as the prior for snapshot $t+1$, thereby accumulating knowledge in a principled manner. This sequential update is order‑insensitive, effectively resists forgetting during evolution, and—through uncertainty quantification(Ahn et al. 2019) (e.g., probabilistic weight distributions)—balances learning and forgetting, avoids catastrophic memory, retains critical knowledge, and gradually discards obsolete information.
Moreover, without constraints, representations of entities and relations may drift across snapshots, exacerbating catastrophic forgetting because the model lacks guidance on how knowledge should evolve while remaining consistent with historical representations. To address this, we propose Continual Clustering, which introduces a contrastive clustering regularizer that keeps entity embeddings in compact clusters during evolution, ensuring that an entity “remembers who it is” and “whom it is related to,” so that semantically similar entities maintain their relative positions over time. Our contributions are as follows:
@@
  • We propose BAKE, a Bayesian‑guided continual knowledge graph embedding approach that formulates CKGE as a sequential Bayesian inference problem, providing theoretical guarantees against catastrophic forgetting.
  • We introduce a continual clustering approach that constrains the evolution of knowledge between KG snapshots, preserving semantic consistency while allowing controlled adaptation to new knowledge.
@@
  • We conduct extensive experiments on multiple CKGE benchmarks and demonstrate that BAKE significantly outperforms existing baseline models in terms of both knowledge preservation and adaptability.
Related Work
Knowledge Graph Embedding
@@
Early KGE models mainly follow the translation paradigm, such as TransE(Bordes et al. 2013) and TransH(Wang et al. 2014) in the real vector space, and their extensions to complex/quaternion spaces, including RotatE(Sun et al. 2019), QuatE(Zhang et al. 2019), HAKE(Zhang et al. 2020), MRME(Li et al. 2025b), and GIE(Cao et al. 2022). These approaches introduced the idea of representing a triple $(h,r,t)$ through simple algebraic operations in a vector space. Since then, both the efficiency and the expressive power of KGE models have improved along multiple dimensions. RecPiece(Liang et al. 2024a) clusters relation centres to select anchors, reducing memory consumption while improving link‑prediction accuracy; RAA‑KGC(Yuan et al. 2025a) pulls queries of a pretrained language model toward relation‑aware neighbour anchors to construct more discriminative embeddings; NFE(Xiao, He, and Cao 2023) treats entities and relations as permutations of random variables, unifying point and probabilistic embeddings under group theory and enabling logical reasoning with uncertainty. Recently, KGE research has begun to leverage diffusion models to recast the task as conditional entity generation. For example, KGDM(Long et al. 2024) trains a diffusion model that outputs the distribution of tail entities conditioned on the head entity and relation, thereby capturing multimodal relational semantics without relying on complicated probabilistic mechanisms.S2DN(Ma et al. 2025) proposes a semantic-aware denoising network and filters out unreliable interactions to enhance the structural reliability of target links.
Continual Knowledge Graph Embedding
@@
As a knowledge graph (KG) evolves, embedding models must learn new facts while avoiding catastrophic forgetting of previously acquired knowledge. Consequently, research on continual KGE (CKGE) has advanced rapidly in recent years. IncDE(Liu et al. 2024a) orders new triples hierarchically and employs layer‑wise incremental distillation to preserve old knowledge; FastKGE(Liu et al. 2024b) accelerates model updates with an adaptive low‑rank adapter (IncLoRA), substantially shortening training time while isolating parameters for new snapshots; LKGE(Cui et al. 2023) proposes a masked auto‑encoder combined with transfer regularisation to extend embeddings across multiple snapshots, enabling lifelong learning; ETT‑CKGE(Zhu et al. 2025) replaces costly node scoring with lightweight task‑driven labels and aligns CKGs through simple matrix operations; CLKGE(Cao 2024) couples continual learning with an energy‑based manifold so that new and old knowledge reinforce each other, whereas CMKGE(Song et al. 2024) introduces a biologically inspired dual inhibition excitation mask to balance plasticity and stability; FMR(Zhu et al. 2024) rotates the parameter space based on Fisher information and assigns flexible regularisation to emulate the refinement of human memory and further mitigate forgetting.
Preliminaries
Continual knowledge graph
@@
A Continual Knowledge Graph (CKG) is represented as a series of snapshots $\mathcal{G}_{c}={\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{N}}$, where each snapshot $\mathcal{S}_{i}={\mathcal{E}_{i},\mathcal{R}_{i},\mathcal{T}_{i}}$ describes the state of the knowledge graph at time $i$. Each triple in $\mathcal{S}_{i}$ is represented as $(e_{h},r,e_{t})\in\mathcal{T}_{i}$, where $e_{h}$ and $e_{t}$ are the head and tail entities drawn from $\mathcal{E}_{i}$, and $r$ is a relation from the relation set $\mathcal{R}_{i}$. This study focuses on scenarios where relations remain constant over time (i.e., $\mathcal{R}_{i}=\mathcal{R}$ for all $i$), but entities expand gradually. The new entities and triples at time $i$ are denoted by $\Delta\mathcal{E}_{i}=\mathcal{E}i\setminus\mathcal{E}{i-1}$ and $\Delta\mathcal{T}_{i}=\mathcal{T}i\setminus\mathcal{T}{i-1}$, respectively. This setting aligns with typical real-world scenarios and is applied in current continual knowledge graph embedding methods.
Continual Knowledge Graph Embedding
[…]
Continual learning and Bayesian neural networks
@@
Exact inference is usually intractable for deep networks, so an independent Gaussian (mean‑field) approximation is commonly adopted (Lee et al. 2024; Ahn et al. 2019): $q_{t}(\boldsymbol{\omega})=\prod_{i=1}^{P}\mathcal{N}\!\bigl{(}\omega_{i};\mu_{t,i},\sigma_{t,i}^{2}\bigr{)}$, with $\theta_{t}=(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ taken as the learnable variational parameters. The independence assumption greatly simplifies the KL‐divergence computation and the reparameterization gradient estimator, enabling iterative posterior updates. In practice, the reparameterization trick $\omega_{i}=\mu_{t,i}+\sigma_{t,i}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ is often employed to estimate gradients.
Methodology
Framework Overview
@@
The overall framework of BAKE is shown in Figure 2. Its core idea is to formalize the continual knowledge graph embedding (CKGE) problem as a serialized Bayesian inference process. Given a knowledge graph sequence $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}$ that evolves over time, the model receives a set of newly added triples $\Delta\mathcal{T}_{t}$ at each time snapshot $t$. BAKE works together through the following three modules: First, Bayesian-guided knowledge evolution uses the parameter posterior distribution of the previous snapshot $t-1$ as a priori, integrates new knowledge and quantifies uncertainty through Bayesian updating; second, the continual clustering module introduces fairness regularization to maintain the semantic consistency of the embedding space; finally, the joint optimization objective integrates the above modules to achieve a balance between new knowledge learning and old knowledge retention. This framework aims to continually accumulate knowledge and alleviate catastrophic forgetting guided by the Bayesian rule.
Figure 2: An overview of our proposed BAKE framework.
Bayesian‐Guided Knowledge Evolution Learning

### v2

@@
Abstract.
As social media and the World Wide Web become hubs for information dissemination, effectively organizing and understanding the vast amounts of dynamically evolving Web content is crucial. Knowledge graphs (KGs) provide a powerful framework for structuring this information. However, the rapid emergence of new hot topics, user relationships, and events in social media renders traditional static knowledge graph embedding (KGE) models rapidly outdated. Continual Knowledge Graph Embedding (CKGE) aims to address this issue, but existing methods commonly suffer from catastrophic forgetting, whereby older, but still valuable, information is lost when learning new knowledge (such as new memes or trending events). This means the model cannot effectively learn the evolution of the data. We propose a novel CKGE framework, BAKE. Unlike existing methods, BAKE formulates CKGE as a sequential Bayesian inference problem and utilizes the Bayesian posterior update principle as a natural continual learning strategy. This principle is insensitive to data order and provides theoretical guarantees to preserve prior knowledge as much as possible. Specifically, we treat each batch of new data as a Bayesian update to the model’s prior. By maintaining the posterior distribution, the model effectively preserves earlier knowledge even as it evolves over multiple snapshots. Furthermore, to constrain the evolution of knowledge across snapshots, we introduce a continual clustering method that maintains the compact cluster structure of entity embeddings through a regularization term, ensuring semantic consistency while allowing controlled adaptation to new knowledge. We conduct extensive experiments on multiple CKGE benchmarks, which demonstrate that BAKE achieves the top performance in the vast majority of cases compared to existing approaches.
Knowledge Graph, Clustering, Continual Learning
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXcopyright: noneccs: Computing methodologies Artificial intelligenceccs: Information systems Data mining
1. Introduction
Figure 1. TKG uses a timestamp quadruple to explicitly annotate temporal information, while CKG decomposes the knowledge graph into time-unmarked evolutionary snapshots, focusing on the continuous updating and migration of knowledge.
@@ 1. Introduction
Knowledge graphs (KGs)(Liang et al., 2024b)(Ji et al., 2021) have become indispensable tools for representing structured information, supporting a wide range of knowledge driven applications such as LLM based question answering(Pan et al., 2024)(Xu et al., 2025) and recommender systems(Zhao et al., 2024)(Yuan et al., 2025b). Knowledge graph embedding (KGE) aims to encode entities and relations in a KG as continuous vectors, and the quality of these embeddings directly affects downstream performance. Classic KGE models (Bordes et al., 2013)(Sun et al., 2019) perform excellently on static KGs, yet real world graphs usually evolve continuously. For example, the YAGO knowledge base evolved from YAGO3(Mahdisoltani et al., 2013) to YAGO4.5(Suchanek et al., 2024), with the number of entities soaring from about 4.6 million to 49 million and facts (triples) from 40 million to 132 million. For such continually evolving knowledge graphs (CKGs), existing KGE models must retrain the entire graph after each update—a computationally prohibitive and impractical strategy for large scale graphs like YAGO. Hence, this work focuses on CKGE models capable of learning CKGs efficiently.
Recent CKGE research focuses on optimizing training structure, parameter efficiency, and regularization and masking mechanisms. In terms of structural optimization, IncDE (Liu et al., 2024a) employs a layered strategy and incremental distillation mechanism to fully leverage graph structure information. In terms of parameter efficiency, FastKGE (Liu et al., 2024b) reduces training parameters through incremental low-rank adapters, and ETT-CKGE (Zhu et al., 2025) uses task-driven tokens for efficient knowledge transfer. In terms of regularization and masking mechanisms, existing methods include lifelong learning based on masked autoencoders (Cui et al., 2023), alignment of old and new knowledge based on energy models (Cao, 2024), biologically inspired dual masking mechanisms (Song et al., 2024), and flexible regularization based on the Fisher information matrix (Zhu et al., 2024). However, these methods inherently focus on passively ”suppressing” forgetting through heuristic regularization terms or complex masking mechanisms, limiting the drastic changes in model parameters when learning new knowledge. While they mitigate catastrophic forgetting to some extent, they fail to provide a fundamental, principled framework for actively guiding knowledge evolution. In other words, they lack a foundational principle for clearly modeling how knowledge accumulates over time, instead focusing on preventing parameter corruption rather than guiding parameter evolution correctly.
@@
As shown in Figure 1, CKGs differ from temporal KGs(Wang et al., 2023; Ding et al., 2025). While the latter model explicitly models temporal dependencies through timestamp relationships, CKGs focus on incremental updates to the graph structure itself, such as the emergence of new entities or relationships over time(Cui et al., 2023). We keenly observe that due to the lack of direct temporal semantics in persistent knowledge graphs, their evolution is perfectly aligned with the sequential update mechanism of the Bayesian framework. This addresses the shortcomings of existing methods in actively guiding knowledge evolution. Following the Bayesian principle (Lee et al., 2024b; Bonnet et al., 2025), we use the posterior distribution of the KG at snapshot $t$ as the prior for snapshot $t+1$, thereby gradually accumulating knowledge in a principled manner. This sequential update mechanism is insensitive to the order of knowledge and can effectively resist the forgetting of previous knowledge during the process of knowledge evolution. At the same time, it balances learning and forgetting through uncertainty quantification(Ahn et al., 2019) (such as the probability distribution of weights), avoids catastrophic memory problems, and ensures that key knowledge is retained while outdated information is gradually released.
Furthermore, suppressing forgetting at the parameter level alone is insufficient. We identified another key challenge as semantic drift at the representation level: the evolution of knowledge representations between KG snapshots needs to be constrained, otherwise entity and relation embeddings in the latent space will drift unordered from snapshot to snapshot. This uncontrolled drift directly destroys the geometric structure of the embedding space, exacerbating catastrophic forgetting because the model cannot maintain the relative semantic relationships between entities, rendering distance- or angle-based scoring functions ineffective. To address this issue, we proposed a continual clustering method. This method acts as a semantic regularizer, constraining entity embeddings to maintain a compact cluster structure during their evolution through contrastive clustering. This ensures that entities remember not only who they are (intra-cluster compactness) but also who they are related to (inter-cluster separability), better ensuring that semantically similar entities maintain their relative positions in the embedding space over time, providing the model with clear guidance on how knowledge should be historically consistent. Our contributions are as follows:
@@
  • We propose the BAKE model, a novel Bayesian-guided CKGE model. This model treats CKGE as a sequential Bayesian inference problem, offering a novel approach to mitigating catastrophic forgetting.
  • We propose a continual clustering method for CKGE that constrains the evolution of knowledge between KG snapshots, maintaining semantic consistency while allowing controlled adaptation to new knowledge. This enables the model to remember not only the identity of entities (intra-cluster compactness) but also their relationships (inter-cluster separability).
@@
  • We conduct extensive experiments on eight CKGE datasets. Across various experimental results, the BAKE model significantly outperforms existing state-of-the-art methods in both knowledge preservation and adaptability.
2. Related Work
2.1. Knowledge Graph Embedding
@@
Early KGE models mainly follow the translation paradigm, such as TransE(Bordes et al., 2013) and TransH(Wang et al., 2014) in the real vector space, and their extensions to complex/quaternion spaces, including RotatE(Sun et al., 2019), QuatE(Zhang et al., 2019), HAKE(Zhang et al., 2020), SATKGC(Ko et al., 2025), MRME(Li et al., 2025a), SymAgent(Liu et al., 2025), HySAE(Li et al., 2025c), and GIE(Cao et al., 2022). These approaches introduced the idea of representing a triple $(h,r,t)$ through simple algebraic operations in a vector space. Since then, both the efficiency and the expressive power of KGE models have improved along multiple dimensions. RecPiece (Liang et al., 2024a) clusters relation centres to select anchors, reducing memory consumption while improving link prediction accuracy; RAAKGC(Yuan et al., 2025a) pulls queries of a pretrained language model toward relation aware neighbour anchors to construct more discriminative embeddings; FedE(Zhou et al., 2024) pioneers a poisoning attack on federated KG embedding, using inferred relations to craft stealthy updates that induce specific errors without degrading main task performance. Recently, KGE research has begun to leverage diffusion models to recast the task as conditional entity generation. For example, FDM(Long et al., 2024) trains a diffusion model that outputs the distribution of tail entities conditioned on the head entity and relation, thereby capturing multimodal relational semantics without relying on complicated probabilistic mechanisms. S2DN(Ma et al., 2025) proposes a semantic-aware denoising network and filters out unreliable interactions to enhance the structural reliability of target links.
2.2. Continual Knowledge Graph Embedding
@@
As knowledge graphs (KGs) evolve, embedding models must learn new facts while avoiding catastrophic forgetting of previously acquired knowledge. Research on continual knowledge graph embedding (CKGE) has progressed notably in recent years. IncDE (Liu et al., 2024a) orders new triples hierarchically and employs layer-wise incremental distillation to preserve old knowledge; FastKGE (Liu et al., 2024b) accelerates updates with an adaptive low-rank adapter (IncLoRA), substantially shortening training time while isolating parameters for new snapshots; LKGE (Cui et al., 2023) combines a masked autoencoder with transfer regularization to extend embeddings across multiple snapshots; ETT-CKGE (Zhu et al., 2025) replaces costly node scoring with lightweight task-driven labels and aligns KG snapshots through simple matrix operations; CLKGE (Cao, 2024) couples continual learning with an energy-based manifold so that new and old knowledge reinforce each other; CMKGE (Song et al., 2024) introduces a biologically inspired dual inhibition–excitation mask to balance plasticity and stability; FMR (Zhu et al., 2024) rotates the parameter space based on Fisher information and assigns flexible regularization to further mitigate forgetting; ERPP (Yang et al., 2025) transfers relational paths across snapshots to shift from catastrophic forgetting to knowledge accumulation; and SAGE (Li et al., 2025b) adaptively expands embedding dimensions and uses dynamic distillation for CKGE.
3. Preliminaries
3.1. Continual knowledge graph
@@
A Continual Knowledge Graph (CKG) is represented as a series of snapshots $\mathcal{G}_{c}={\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{N}}$, where each snapshot $\mathcal{S}_{i}={\mathcal{E}_{i},\mathcal{R}_{i},\mathcal{T}_{i}}$ describes the state of the knowledge graph at time $i$. Each triple in $\mathcal{S}_{i}$ is represented as $(e_{h},r,e_{t})\in\mathcal{T}_{i}$, where $e_{h}$ and $e_{t}$ are the head and tail entities drawn from $\mathcal{E}_{i}$, and $r$ is a relation from the relation set $\mathcal{R}_{i}$. This study focuses on scenarios where relations remain constant over time (i.e., $\mathcal{R}_{i}=\mathcal{R}$ for all $i$), but entities expand gradually. The new entities and triples at time $i$ are denoted by $\Delta\mathcal{E}_{i}=\mathcal{E}_{i}-\mathcal{E}_{i-1}$ and $\Delta\mathcal{T}_{i}=\mathcal{T}_{i}-\mathcal{T}_{i-1}$, respectively. This setting aligns with typical real-world scenarios and is applied in current continual knowledge graph embedding methods.
3.2. Continual Knowledge Graph Embedding
The goal of Knowledge Graph Embedding is to embed entities and relations into a low-dimensional vector space while preserving their semantic meanings. For a CKG $\mathcal{G}_{c}={\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{i}}$, CKGE progressively builds an embedding model $\mathcal{M}_{i}$ as new snapshots $\mathcal{S}_{i}$ arrive. It leverages prior embeddings from snapshots $\mathcal{S}_{0},\dots,\mathcal{S}_{i-1}$}, optimizing training efficiency and overall performance. Completely retraining $\mathcal{M}_{i}$ from scratch on all triples in $\mathcal{S}_{i}$ is a basic yet inefficient choice unsuitable for practical scenarios. Current CKGE methods adapt the existing model $\mathcal{M}_{i-1}$ to new data $\mathcal{M}_{i}$ by creating embeddings for new entities $\Delta\mathcal{E}_{i}$ and refining embeddings for existing relations $\mathcal{R}$ and prior entities $\mathcal{E}_{i-1}$ when receiving new triples $\Delta\mathcal{T}_{i}$. CKGE is evaluated through continual link prediction: given an incomplete triple such as $(h,r,?)$, the model scores each candidate tail entity and outputs the highest-ranked one, maintaining accuracy while learning new knowledge and remembering past knowledge.
3.3. Continual learning and Bayesian neural networks
@@
Exact inference is usually intractable for deep networks, so an independent Gaussian (mean field) approximation is commonly adopted (Lee et al., 2024b; Ahn et al., 2019): $q_{t}(\boldsymbol{\omega})=\prod_{i=1}^{P}\mathcal{N}\!\bigl(\omega_{i};\mu_{t,i},\sigma_{t,i}^{2}\bigr)$, with $\theta_{t}=(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ taken as the learnable variational parameters. The independence assumption greatly simplifies the KL divergence computation and the reparameterization gradient estimator, enabling iterative posterior updates. In practice, the reparameterization trick $\omega_{i}=\mu_{t,i}+\sigma_{t,i}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ is often employed to estimate gradients.
4. Methodology
4.1. Framework Overview
@@
The overall framework of BAKE is shown in Figure 2. Its core idea is to formalize the continual knowledge graph embedding (CKGE) problem as a serialized Bayesian inference process. Given a knowledge graph sequence $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}$ that evolves over time, the model receives a set of newly added triples $\Delta\mathcal{T}_{t}$ at each time snapshot $t$. BAKE works together through the following three modules: First, Bayesian-guided knowledge evolution uses the parameter posterior distribution of the previous snapshot t1t-1 as a priori, integrates new knowledge and quantifies uncertainty through Bayesian updating; second, the continual clustering module introduces fairness regularization to maintain the semantic consistency of the embedding space; finally, the joint optimization objective integrates the above modules to achieve a balance between new knowledge learning and old knowledge retention. This framework aims to continually accumulate knowledge and alleviate catastrophic forgetting guided by the Bayesian rule.
Figure 2. An overview of our proposed BAKE framework.
4.2. Bayesian Guided Knowledge Evolution Learning